The perimeter of a square is 36 centimeters. What is the area of the square?
Answer by mananth(16946)

The perimeter of a square is 36 centimeters. What is the area of the square?
Perimeter is = summ of all side.
Since it is a square the sides are equal
36/4 = 9. side = 9 cms.
Area of square = side * side
= 9* 9 = 81 square cms.